Calculate the maximum deposit outflow a bank can sustain without altering its balance sheet given the amount of checkable deposits, the required reserve ratio, and the amount held in reserves.

Solve for x

reserves= 40,000

required reserve ratio= .20

checkable deposits= 100,000

please show how to solve. thanks!!
Answer #1

Given that,

Reserves (R) = 40,000

Required reserve ratio = .20

Checkable Deposits (D) = 100,000

Hence, the maximum deposit outflow it can sustain without altering its balance sheet is

(R - %D)/ (1-%) = {40,000 - (.2*100,000) } / (1 - 0.2) = (40,000 - 20,000) / (0.8) = 20,000/0.8 = 25,000